OpenStudy (cristiancha5162):

Consider the following pair of equations: x + y = −2 y = 2x + 10 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−4, 2) (4, 2) (−2, 4) (2, 4)

OpenStudy (cristiancha5162):

@AnitaS.

OpenStudy (cristiancha5162):

@Qwertty123

OpenStudy (cristiancha5162):

i think its b

OpenStudy (anonymous):

Substitute the second equation into the first to get an expressiin solely in term of x. Solve the resulting equation to get a value for x. If you still don't have enough info, take this value, put it into one of the original equations. This will give you a value for y.

OpenStudy (cristiancha5162):

is it b

OpenStudy (g3n3sis08):

no its A

OpenStudy (cristiancha5162):

why

OpenStudy (anonymous):

Why do you think it's b? Alternatively, take each pair of values and substitute them into both equations. Only the pair which evaluates as true for both equations at once can be the answer.

OpenStudy (cristiancha5162):

so its a

OpenStudy (g3n3sis08):

y=2x+10 x+y=-2 y=-2-x -2-x=2x+10 0=3x+12 x=-4 y=-2-(-4) y=-2+4 y=2 (-4,2)

OpenStudy (cristiancha5162):

tthnx

OpenStudy (g3n3sis08):

welcome :)

OpenStudy (g3n3sis08):

welcome :)